OpenStudy (anonymous):

Hailey's personal goal is to raise at least $150. Solve the inequality 20 + 5s = 150 to determine the number of laps Hailey needs to swim, s, to meet her goal. Which of the following most accurately describes the solution?

OpenStudy (anonymous):

can i get the answer?! LOL

OpenStudy (anonymous):

Subtract 20 from 150 which gives you 130. remaining is 5s=130 divide 5s-5=130/5 which gives you 26. so s=26 laps. does this help?

OpenStudy (anonymous):

yes!! hehe ;)

OpenStudy (anonymous):

:)